OpenStudy (anonymous):

a box contains 2 defective lights inadvertenly mixed with 8 nondefective lights. If the lights are selected one at a time without replacement and tested until both defective lights are found, what is the probability that both defective lights will be found after exactly 3 trials

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

so take a general case (D)(D)(ND) where D= defective , ND = not defective

OpenStudy (anonymous):

the probability of this one event happening is P(D, D, ND) = 2/10 x 1/9 ( we stop after they found the two bad ones )

OpenStudy (anonymous):

P(D,ND,D) is the only other case to consider

OpenStudy (anonymous):

a couple ways to do this but easiest might be this. there are 3 possibilities for getting exactly 2 defective in 3 tries: ddn dnd ndd where d is defective and n is not. each of these is given by the probability \[\frac{2}{8}\times \frac{1}{7} \times\frac{6}{6}\] so answer is \[3\times \frac{2}{8}\times \frac{1}{7}=\frac{3}{28}\]

OpenStudy (anonymous):

doesnt quite work

OpenStudy (anonymous):

that way

OpenStudy (anonymous):

messed that all up didn't i?

OpenStudy (anonymous):

I think theres something wrong, ohh might be right, probabilities is not my very best

OpenStudy (anonymous):

right idea though. two out of eight. \[\frac{2}{10}\times \frac{1}{9}\times \frac{8}{8}\times 3\]

OpenStudy (anonymous):

unless "exactly three:" means on the third trial in which case it is dnd ndd but not ddn

OpenStudy (anonymous):

yeh thats was what I was thinking lol

OpenStudy (anonymous):

it says they stop after they found duds

OpenStudy (anonymous):

not clear is it?

OpenStudy (anonymous):

so u cant include them

OpenStudy (anonymous):

well "exactly 3" might mean "it takes three trials" or it might mean "in three trials we have two defective" not sure which

OpenStudy (anonymous):

if it means "in three trials we have two defective" i think my answer is right

OpenStudy (anonymous):

if it means "it takes exactly three trials before we know" we have only two possibilities DND NDD

OpenStudy (anonymous):

second case would be \[2\times \frac{2}{10}\times \frac{8}{9}\times \frac{1}{8} \]